ADVANCED MATERIAL Calculating the solubility of an lonic compound when a common lon is present Calculate the solubility at 25 Ksp = 3.45 • 10^-11
Answer #1

Cafe – Ca 2+ + 2 + U S 25. Ksp = [Ca 2+] [F-] 2 3.45 X 10 11 us3 so 2:05x154 mol 3 - 2005X10 4 x (19.07) = i S= 0.0163/ 0.016